# Coderivative-Based Newton Methods in Structured Nonconvex and Nonsmooth Optimization

### v2

@@ 4 Local Convergence Analysis of Coderivative-Based Newton Method for Structured Nonconvex Optimization
Let $\bar{x}$ be a critical point of the problem (4.1) for which $\nabla f+\partial g$ is metrically regular around $(\bar{x},0)$ and $\partial g$ is semismooth at $(\bar{x},-\nabla f(\bar{x}))$. Suppose further that $g$ is continuously prox-regular at $\bar{x}$ for $-\nabla f(\bar{x})$ with modulus $r>0$. Then for $\lambda\in(0,r^{-1})$, there exists a neighborhood $U$ of $\bar{x}$ such that for all $x^{0}\in U$ we have that Algorithm 2 either stops after finitely many iterations, or produces a sequence $\{x^{k}\}$ that converges $Q$-superlinearly to $\bar{x}$.
[Proof.] The limiting subdifferential sum rule from [39, Proposition 1.107] tells us that
@@
$\partial_{B}f(\bar{x}):=\big{\{}H\in{\rm I\!R}^{m\times n}\;\big{|}\;\exists\ x_{k}\to\bar{x}\text{ such that }f\text{ is differentiable at }x_{k}\;\text{ and }\;\nabla f(x_{k})\to H\big{\}}.$ (4.17)
@@
Let $\varphi:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ be defined in (4.1). For $\bar{x}\in\mbox{\rm dom}\,\varphi$ and $\bar{v}\in\partial\varphi(\bar{x})$, suppose that $g:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ is $r$-level prox-regular at $\bar{x}\in\mbox{\rm dom}\,\varphi$ for $\bar{v}-\nabla f(\bar{x})$. If the subgradient mapping $\partial\varphi$ is metrically regular around $(\bar{x},\bar{v})$, then for any $\lambda\in(0,1/r)$ and $A\in\partial_{B}(\text{\rm Prox}_{\lambda g})(\bar{x}+\lambda(\bar{v}-\nabla f(\bar{x})))$, we have the assertions:
@@
$-\lambda A\bar{v}=(I-A+\lambda A\nabla^{2}f(\bar{x}))d,$ (4.19)
@@
[Proof.] To verify (i), pick $\lambda\in(0,1/r)$ and $A\in\partial_{B}(\text{\rm Prox}_{\lambda g})(\bar{x}+\lambda(\bar{v}-\nabla f(\bar{x})))$. Take any $u\in{\rm I\!R}^{n}$ satisfying
$(I-A+\lambda A\nabla^{2}f(\bar{x}))u=0$ (4.20)
and show that $u=0$. To proceed, rewrite (4.20) in the form
$-\nabla^{2}f(\bar{x})u=\frac{1}{\lambda}(I-A)(u-\lambda\nabla^{2}f(\bar{x})u).$ (4.21)
@@
$-\nabla^{2}f(\bar{x})u\in\partial^{2}e_{\lambda}g(\bar{x}+\lambda(\bar{v}-\nabla f(\bar{x})))(u-\lambda\nabla^{2}f(\bar{x})u).$
@@
$-\nabla^{2}f(\bar{x})u\in\partial^{2}g(\bar{x},\bar{v}-\nabla f(\bar{x}))(u),$
which is equivalent to $0\in\partial^{2}\varphi(\bar{x},\bar{v})(u)$ by the second-order sum rule from [39, Proposition 1.121]. Since $\partial\varphi$ is metrically regular around $(\bar{x},\bar{v})$, it follows from the Mordukhovich criterion (2.8) that $u=0$ as claimed in (i).
@@
$-\bar{v}-\nabla^{2}f(\bar{x})d=\frac{1}{\lambda}(I-A)(d+\lambda(-\bar{v}-\nabla^{2}f(\bar{x})d)),$
@@
$-\bar{v}-\nabla^{2}f(\bar{x})d\in\partial^{2}e_{\lambda}g(\bar{x}+\lambda(\bar{v}-\nabla f(\bar{x})))(d+\lambda(-\bar{v}-\nabla^{2}f(\bar{x})d)).$
@@
$-\bar{v}-\nabla^{2}f(\bar{x})d\in\partial^{2}g(\bar{x},\bar{v}-\nabla f(\bar{x}))(d),$
and thus $d$ satisfies (3.4) by the second-order sum rule from [39, Proposition 1.121].
@@ 5 Generalized Line Search Proximal Gradient Method
Let $\varphi$ be given in (4.1). The forward-backward envelope ((((FBE)))) of $\varphi$ with the parameter value $\lambda>0$ is given by
@@
Recall from [24, Proposition 7.1] that under fairly mild assumptions, the FBE $\varphi_{\lambda}$ is ${\cal C}^{1}$-smooth with the explicit calculation of its gradient and the preservation of stationary points of $\varphi$.
@@
Let $f$ be of class $\mathcal{C}^{2}$ around $\bar{x}$ where $0\in\partial\varphi(\bar{x})$, and let $g$ be prox-bounded and continuously prox-regular at $\bar{x}$ for $-\nabla f(\bar{x})$ with modulus $r>0$. Then for all $\lambda\in(0,r^{-1})$, there exists a neighborhood $U$ of $\bar{x}$ on which the FBE $\varphi_{\lambda}$ is continuously differentiable and its gradient mapping is calculated by
@@
$\nabla\varphi_{\lambda}(x)=0\iff 0\in\partial\varphi(x).$ (5.7)
@@
(ii) The FBE $\varphi_{\lambda}$ is of class $\mathcal{C}^{1,1}$ around $\bar{x}$ if the mapping $\nabla^{2}f$ is locally Lipschitzian around $\bar{x}$. Indeed, Lemma 2.5 tells us that the Moreau envelope $e_{\lambda}g$ is of class $\mathcal{C}^{1,1}$ around $\bar{x}-\lambda\nabla f(\bar{x})$. Combining this with the FBE representation in form (5.5) verifies the claim.
@@
Let $\bar{x}\in{\rm I\!R}^{n}$ be a critical point of $\varphi=f+g$ in (4.1). Assume that $\nabla^{2}f$ is strictly differentiable at $\bar{x}$ and that $g\colon{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ is prox-regular at $\bar{x}$ for $-\nabla f(\bar{x})$ with modulus $r>0$. For any $\lambda\in(0,r^{-1})$ such that $Q(\bar{x}):=I-\lambda\nabla^{2}f(\bar{x})$ is nonsingular, we have the FBE generalized Hessian representation
@@
provided that $g$ is subdifferentially continuous at $\bar{x}$ for $-\nabla f(\bar{x})$.
@@
It follows from the differentiability of $\nabla^{2}f$ at $\bar{x}$ that $\psi$ is twice differentiable there. We have by (5.5) that $\varphi_{\lambda}=\psi+h$ and thus apply to this sum the second-order sum rule from [39, Proposition 1.121] to get
@@
It follows from Proposition 2.5 that $h$ is continuously differentiable at $\bar{x}$. By using the second-order chain rule from [39, Theorem 1.127], we have the representation
@@
Since $\bar{x}$ is a critical point of $\varphi$, it follows from (2.19) that
@@
Applying finally the second-order subdifferential sum rule from [39, Proposition 1.121] gives us
@@
In the setting of Lemma 5.4, suppose that $g$ is subdifferentially continuous at $\bar{x}$ for $-\nabla f(\bar{x})$, and let $\lambda\in(0,r^{-1})$ be such that $Q(\bar{x})=I-\lambda\nabla^{2}f(\bar{x})$ is positive-definite. If $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ in (4.1), then FBE (5.3) is strongly convex around $\bar{x}$.
[Proof.] It follows from the assumptions that $\varphi$ is continuously prox-regular at $\bar{x}$ for 00. Combining this with the pointbased second-order characterization of tilt stability from [49, Theorem 1.3] gives us the positive-definiteness of $\partial^{2}\varphi(\bar{x},0)$. Note that the strict differentiability of $\nabla^{2}f$ yields the local Lipschitz continuity of $\nabla^{2}f$ around $\bar{x}$. Therefore, $\varphi_{\lambda}$ is of class $\mathcal{C}^{1,1}$ around $\bar{x}$ due to Remark 5.3(ii). By [7, Proposition 4.6], we verify the strong convexity of $\varphi_{\lambda}$ around $\bar{x}$ via showing that $\partial^{2}\varphi_{\lambda}(\bar{x})$ is positive-definite. To furnish this, pick $w\in{\rm I\!R}^{n}\setminus\{0\}$ and $z\in\partial^{2}\varphi_{\lambda}(\bar{x})(w)$. Then we deduce from (5.10) that $z\in\partial^{2}\varphi(\bar{x},0)(w-\lambda Q(\bar{x})^{-1}z)$, which ensures that $z\neq 0$ by the positive-definiteness of $\partial^{2}\varphi(\bar{x},0)$. This also implies that $\langle z,w-\lambda Q(\bar{x})^{-1}z\rangle\geq 0$, and hence
$\langle z,w\rangle\geq\lambda\langle Q(\bar{x})^{-1}z,z\rangle>0$ (5.18)
due to the positive-definiteness of $Q(\bar{x})^{-1}$. In this way, which verifies the claim.
@@
1:$x^{0}\in{\rm I\!R}^{n}$, $\lambda\in(0,\text{\rm min}\{1/L_{f},\lambda_{g}\})$, $\sigma\in\left(0,\frac{\lambda(1-\lambda L_{f})}{2(1+\lambda L_{f})^{2}}\right)$, $\beta\in(0,1)$
@@
$\widehat{x}\in\text{\rm Prox}_{\lambda g}(x-\lambda\nabla f(x))\;\mbox{ and }\>\widehat{v}:=\nabla f(\widehat{x})-\nabla f(x)+\frac{1}{\lambda}(x-\widehat{x}).$ (5.19)
@@
Consequently, both sequences $\{\widehat{v}^{k}\}$ and $\{\widehat{x}^{k}-x^{k}\}$ converge to 00 as $k\to\infty$. Finally, any accumulation point of $\{x^{k}\}$ is a critical point of (4.1).
[Proof.] It follows from Lemma 5.6 that Algorithm 3 is well-defined and exhibits the descent property $\varphi_{\lambda}(x^{k+1})<\varphi_{\lambda}(x^{k})$. Arguing similarly to the proof of inequalities (5.22) and (5.23) brings us to (5.27). By [58, Theorem 4.4], we have $\inf\varphi_{\lambda}=\inf\varphi>-\infty$. Remembering that the sequence $\{\varphi_{\lambda}(x^{k})\}$ is monotonically decreasing yields the convergence $\varphi_{\lambda}(x^{k})\to\inf\varphi_{\lambda}$ as $k\to\infty$. Moreover, we have the estimates
@@
which justify (5.28) and ensures that both sequences $\{\widehat{v}^{k}\}$ and $\{\widehat{x}^{k}-x^{k}\}$ converge to 00 as $k\to\infty$.
@@
$0=\lim_{k\to\infty,k\in J}\|\widehat{v}^{k}\|=\|\nabla f(\widehat{x})+\lambda^{-1}(\bar{x}-\lambda\nabla f(\bar{x})-\widehat{x})\|,$
which implies in turn that $0\in\mathcal{F}_{\lambda}^{\text{\rm nor}}(\bar{x}-\lambda\nabla f(\bar{x}))$. Using Proposition 4.5, we conclude that $\bar{x}$ is a critical point of (4.1) and thus complete the proof of the proposition.
@@
[Proof.] It follows from Proposition 5.8 that $\bar{x}$ is a critical point of (4.1) with $\widehat{v}^{k}\to 0$ and $\widehat{x}^{k}-x^{k}\to 0$ as $k\to\infty$. Let a subsequence $\{x^{k}\}_{k\in J}$ converge to $\bar{x}$. For all large $k$, we get from (5.29) that
@@
[…]
@@
Let $\bar{x}$ be a critical point of (4.1). Suppose that:
@@
(iii) $g$ is continuously prox-regular at $\bar{x}$ for $-\nabla f(\bar{x})$ with modulus $r>0$.
@@
In the setting of Proposition 5.9, suppose that the accumulation point $\bar{x}$ satisfies all the conditions in Assumption 2. In Algorithm 3, choose $\lambda\in(0,r^{-1})$ such that $Q(\bar{x})=I-\lambda\nabla^{2}f(\bar{x})$ is positive-definite. Then we have the assertions:
@@
[Proof.] It follows from Proposition 5.8 that $\widehat{v}^{k}\to 0$ with the estimates in (5.27). Invoking Proposition 5.9 ensures that $x^{k}\to\bar{x}$ as $k\to\infty$. By the prox-regularity of $g$ at $\bar{x}$ for $-\nabla f(\bar{x})$, Proposition 4.7 tells us that the mapping $x\mapsto\text{\rm Prox}_{\lambda g}(x-\lambda\nabla f(x))$ is single-valued around $\bar{x}$. Then we deduce from Remark 5.3(i), the criticality of $\bar{x}$, and (5.8) that $\nabla\varphi_{\lambda}(\bar{x})=0$. It follows from Remark 5.3(ii) that $\varphi_{\lambda}$ is of class $\mathcal{C}^{1,1}$ around $\bar{x}$, and thus we get that $\varphi_{\lambda}(x^{k})\to\varphi_{\lambda}(\bar{x})$ and $\|\nabla\varphi_{\lambda}(x^{k})\|\to 0$ as $k\to\infty$. Moreover, Lemma 5.5 implies that $\varphi_{\lambda}$ is strongly convex around $\bar{x}$, and hence there exists $\kappa>0$ such that
$\varphi_{\lambda}(x)\geq\varphi_{\lambda}(\bar{x})+\langle\nabla\varphi_{\lambda}(\bar{x}),x-\bar{x}\rangle+\frac{\kappa}{2}\|x-\bar{x}\|^{2}=\varphi_{\lambda}(\bar{x})+\frac{\kappa}{2}\|x-\bar{x}\|^{2},$ (5.30)

### v3

@@ 4 Local Convergence Analysis of Coderivative-Based Newton Method for Structured Nonconvex Optimization
Let $\bar{x}$ be an M-stationary point of the problem (4.1) for which $\nabla f+\partial g$ is metrically regular around $(\bar{x},0)$ and $\partial g$ is semismooth at $(\bar{x},-\nabla f(\bar{x}))$. Suppose further that $g$ is continuously $r$-level prox-regular at $\bar{x}$ for $\nabla f(\bar{x})$ for some $r>0$, and $\bar{x}$ is a $\bar{\lambda}$-critical point of (4.1) for some $\bar{\lambda}\in(0,1/r]$. Then for $\lambda\in(0,\bar{\lambda})$, there exists a neighborhood $U$ of $\bar{x}$ such that for all $x^{0}\in U$ we have that Algorithm 2 either stops after finitely many iterations, or produces a sequence $\{x^{k}\}$ that converges $Q$-superlinearly to $\bar{x}$.
[Proof.] The limiting subdifferential sum rule from [38, Proposition 1.107] tells us that
@@
[…]
@@
Let $\varphi:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ be defined in (4.1). For $\bar{x}\in\mbox{\rm dom}\,\varphi$ and $\bar{v}\in\partial\varphi(\bar{x})$, suppose that $g:{\rm I\!R}^{n}\to\overline{{\rm I\!R}}$ is $r$-weakly convex for some $r>0$. If the subgradient mapping $\partial\varphi$ is metrically regular around $(\bar{x},\bar{v})$, then for any $\lambda\in(0,1/r)$ and $A\in\partial_{B}(\text{\rm Prox}_{\lambda g})(\bar{x}+\lambda(\bar{v}-\nabla f(\bar{x})))$, we have the assertions:
@@
$\lambda A\bar{v}=(I-A+\lambda A\nabla^{2}f(\bar{x}))d,$ (4.14)
@@
[Proof.] To verify (i), pick $\lambda\in(0,1/r)$ and $A\in\partial_{B}(\text{\rm Prox}_{\lambda g})(\bar{x}+\lambda(\bar{v}-\nabla f(\bar{x})))$. It follows from Proposition 2.5 and Remark 2.6(ii) that $\text{\rm Prox}_{\lambda g}$ is single-valued, Lipschitz continuous, and
$\nabla e_{\lambda}g(x)=\frac{1}{\lambda}(x-\text{\rm Prox}_{\lambda g}(x))\quad\text{for all }\;x\in{\rm I\!R}^{n}.$ (4.15)
Taking any $u\in{\rm I\!R}^{n}$ satisfying
$(I-A+\lambda A\nabla^{2}f(\bar{x}))u=0,$ (4.16)
our goal is to show that $u=0$. To proceed, rewrite (4.16) in the form
$\nabla^{2}f(\bar{x})u=\frac{1}{\lambda}(I-A)(u-\lambda\nabla^{2}f(\bar{x})u).$ (4.17)
@@
$\nabla^{2}f(\bar{x})u\in\partial^{2}e_{\lambda}g(\bar{x}+\lambda(\bar{v}-\nabla f(\bar{x})))(u-\lambda\nabla^{2}f(\bar{x})u).$
@@
$\nabla^{2}f(\bar{x})u\in\partial^{2}g(\bar{x},\bar{v}-\nabla f(\bar{x}))(u),$
which is equivalent to $0\in\partial^{2}\varphi(\bar{x},\bar{v})(u)$ by the second-order sum rule from [38, Proposition 1.121]. Since $\partial\varphi$ is metrically regular around $(\bar{x},\bar{v})$, it follows from the Mordukhovich criterion (2.8) that $u=0$ as claimed in (i).
@@
$\bar{v}-\nabla^{2}f(\bar{x})d=\frac{1}{\lambda}(I-A)(d+\lambda(-\bar{v}-\nabla^{2}f(\bar{x})d)),$
@@
$\bar{v}-\nabla^{2}f(\bar{x})d\in\partial^{2}e_{\lambda}g(\bar{x}+\lambda(\bar{v}-\nabla f(\bar{x})))(d+\lambda(-\bar{v}-\nabla^{2}f(\bar{x})d)).$
@@
$\bar{v}-\nabla^{2}f(\bar{x})d\in\partial^{2}g(\bar{x},\bar{v}-\nabla f(\bar{x}))(d),$
and thus $d$ satisfies (3.4) by the second-order sum rule from [38, Proposition 1.121].
@@ 5 Generalized Line Search Proximal Gradient Method
Let $\varphi$ be given as in (4.1). The forward-backward envelope ((FBE)) of $\varphi$ with the parameter value $\lambda>0$ is given by
@@
It follows from Proposition 2.5 and (5.4) that under fairly mild assumptions, the FBE $\varphi_{\lambda}$ is ${\cal C}^{1}$-smooth with the explicit calculation of its gradient and the preservation of stationary points of $\varphi$.
@@
Let $\bar{x}$ be an M-stationary point of (4.1), and let $g$ be continuously $r$-level prox-regular at $\bar{x}$ for $\bar{v}:=-\nabla f(\bar{x})$ for some $r>0$. Suppose that $\bar{x}$ is a $\bar{\lambda}$-critical point of (4.1) for some $\bar{\lambda}\in(0,1/r).$ Then for all $\lambda\in(0,\bar{\lambda})$, there exists a neighborhood $U$ of $\bar{x}$ on which the FBE $\varphi_{\lambda}$ is continuously differentiable and its gradient mapping is calculated by
@@
$\nabla\varphi_{\lambda}(x)=0\iff 0\in\partial\varphi(x).$
@@
(ii) The FBE $\varphi_{\lambda}$ is of class $\mathcal{C}^{1,1}$ around $\bar{x}$ if the mapping $\nabla^{2}f$ is locally Lipschitzian around $\bar{x}$. Indeed, Proposition 2.5 tells us that the Moreau envelope $e_{\lambda}g$ is of class $\mathcal{C}^{1,1}$ around $\bar{x}-\lambda\nabla f(\bar{x})$. Combining this with the FBE representation in form (5.4) verifies the claim.
@@
Let $\bar{x}$ be an M-stationary point of (4.1), and let $g$ be $r$-level prox-regular at $\bar{x}$ for $\bar{v}:=-\nabla f(\bar{x})$ for some $r>0$. Suppose that $\bar{x}$ is a $\bar{\lambda}$-critical point of (4.1) for some $\bar{\lambda}\in(0,1/r)$, and $\nabla^{2}f$ is strictly differentiable at $\bar{x}$. For any $\lambda\in(0,\bar{\lambda})$ such that $Q(\bar{x}):=I-\lambda\nabla^{2}f(\bar{x})$ is nonsingular, we have the FBE generalized Hessian representation
@@
provided that $g$ is subdifferentially continuous at $\bar{x}$ for $\nabla f(\bar{x})$.
@@
It follows from the differentiability of $\nabla^{2}f$ at $\bar{x}$ that $\psi$ is twice differentiable there. We have by (5.4) that $\varphi_{\lambda}=\psi+h$ and thus apply to this sum the second-order sum rule from [38, Proposition 1.121] to get
@@
It follows from Proposition 2.5 that $h$ is continuously differentiable at $\bar{x}$. By using the second-order chain rule from [38, Theorem 1.127], we have the representation
@@
Since $\bar{x}$ is a $\bar{\lambda}$-critical point of $\varphi$, it is also $\lambda$-critical point due to Remark 4.3. Moreover, it follows from Proposition 2.5 that
@@
Applying finally the second-order subdifferential sum rule from [38, Proposition 1.121] gives us
@@
In the setting of Lemma 5.4, suppose that $g$ is subdifferentially continuous at $\bar{x}$ for $\nabla f(\bar{x})$, and let $\lambda\in(0,\bar{\lambda})$ be such that $Q(\bar{x})=I-\lambda\nabla^{2}f(\bar{x})$ is positive-definite. If $\bar{x}$ is a tilt-stable local minimizer of $\varphi$ in (4.1), then FBE (5.2) is strongly convex around $\bar{x}$.
[Proof.] It follows from the assumptions that $\varphi$ is continuously prox-regular at $\bar{x}$ for 0. Combining this with the pointbased second-order characterization of tilt stability from [48, Theorem 1.3] gives us the positive-definiteness of $\partial^{2}\varphi(\bar{x},0)$. Note that the strict differentiability of $\nabla^{2}f$ yields the local Lipschitz continuity of $\nabla^{2}f$ around $\bar{x}$. Therefore, $\varphi_{\lambda}$ is of class $\mathcal{C}^{1,1}$ around $\bar{x}$ due to Remark 5.3(ii). By [7, Proposition 4.6], we verify the strong convexity of $\varphi_{\lambda}$ around $\bar{x}$ via showing that $\partial^{2}\varphi_{\lambda}(\bar{x})$ is positive-definite. To furnish this, pick $w\in{\rm I\!R}^{n}\setminus\{0\}$ and $z\in\partial^{2}\varphi_{\lambda}(\bar{x})(w)$. Then we deduce from (5.8) that $z\in\partial^{2}\varphi(\bar{x},0)(w-\lambda Q(\bar{x})^{-1}z)$, which ensures that $z\neq 0$ by the positive-definiteness of $\partial^{2}\varphi(\bar{x},0)$. This also implies that $\langle z,w-\lambda Q(\bar{x})^{-1}z\rangle\geq 0$, and hence
$\langle z,w\rangle\geq\lambda\langle Q(\bar{x})^{-1}z,z\rangle>0$
due to the positive-definiteness of $Q(\bar{x})^{-1}$. This verifies the claim.
@@
[…]
@@
[…]
@@
Consequently, both sequences $\{\widehat{v}^{k}\}$ and $\{\widehat{x}^{k}-x^{k}\}$ converge to 0 as $k\to\infty$. Finally, any accumulation point of $\{x^{k}\}$ is a $\lambda$-critical point of (4.1) for all $\lambda\in(0,\text{\rm min}\{1/L_{f},\lambda_{g}\})$.
[Proof.] It follows from Lemma 5.6 that Algorithm 3 is well-defined and exhibits the descent property $\varphi_{\lambda}(x^{k+1})<\varphi_{\lambda}(x^{k})$. Arguing similarly to the proof of inequalities (5.18) and (5.19) brings us to (5.23). By [57, Theorem 4.4], we have $\inf\varphi_{\lambda}=\inf\varphi>$-$\infty$. Remembering that the sequence $\{\varphi_{\lambda}(x^{k})\}$ is monotonically decreasing yields the convergence $\varphi_{\lambda}(x^{k})\to\inf\varphi_{\lambda}$ as $k\to\infty$. Moreover, we have the estimates
@@
which justify (5.24) and ensures that both sequences $\{\widehat{v}^{k}\}$ and $\{\widehat{x}^{k}-x^{k}\}$ converge to 0 as $k\to\infty$.
Suppose further that $\bar{x}$ is an accumulation point of $\{x^{k}\}$ and hence find an infinite set $J\subset{\rm I\!N}$ such that $x^{k}\overset{J}{\rightarrow}\bar{x}$. By Proposition 2.4, we deduce from the inclusion $\widehat{x}^{k}\in\text{\rm Prox}_{\lambda g}(x^{k}-\lambda\nabla f(x^{k}))$ that the sequence $\{\widehat{x}^{k}\}_{k\in J}$ is bounded and that all its accumulation points lie in $\text{\rm Prox}_{\lambda g}(\bar{x}-\lambda\nabla f(\bar{x}))$. Without loss of generality, assume that $\widehat{x}^{k}\overset{J}{\rightarrow}\widehat{x}\in\text{\rm Prox}_{\lambda g}(\bar{x}-\lambda\nabla f(\bar{x}))$. By $\widehat{v}^{k}\to 0$, it yields
@@
which implies in turn that $0\in\mathcal{F}_{\lambda}^{\text{\rm nor}}(\bar{x}-\lambda\nabla f(\bar{x}))$. Using Proposition 4.5, we conclude that $\bar{x}$ is a $\lambda$-critical point of (4.1) and thus complete the proof of the proposition.
@@
[Proof.] It follows from Proposition 5.8 that $\bar{x}$ is a $\lambda$-critical point of (4.1) with $\widehat{v}^{k}\to 0$ and $\widehat{x}^{k}-x^{k}\to 0$ as $k\to\infty$. Let a subsequence $\{x^{k}\}_{k\in J}$ converge to $\bar{x}$. For all large $k$, we get from (5.25) that
@@
which implies that $\|x^{k+1}-x^{k}\|\overset{J}{\rightarrow}0$. Then we deduce from Ostrowski’s condition [11, Proposition 8.3.10]that the entire sequence $\{x^{k}\}$ converges to $\bar{x}$ as $k\to\infty$.
@@
Let $\bar{x}$ be an M-stationary point of (4.1). Suppose that:
@@
(iii) $g$ is continuously $r$-level prox-regular at $\bar{x}$ for $\nabla f(\bar{x})$ for some $r>0$.
@@
In the setting of Proposition 5.9, suppose that the accumulation point $\bar{x}$ satisfies all the conditions in Assumption 2. In Algorithm 3, choose $\lambda\in(0,1/r)$ such that $Q(\bar{x})=I-\lambda\nabla^{2}f(\bar{x})$ is positive-definite. Then we have the assertions:
@@
[Proof.] It follows from Proposition 5.8 that $\widehat{v}^{k}\to 0$ with the estimates in (5.23). Invoking Proposition 5.9 ensures that $x^{k}\to\bar{x}$ as $k\to\infty$, and $\bar{x}$ is a $\lambda$-critical point of (4.1) for all $\lambda\in(0,\text{\rm min}\{1/L_{f},\lambda_{g}\})$. Fix $\lambda$ such that $Q(\bar{x})=I-\lambda\nabla^{2}f(\bar{x})$ is positive-definite. By the $r$-level prox-regularity of $g$ at $\bar{x}$ for $\nabla f(\bar{x})$, Proposition 4.7 tells us that the mapping $x\mapsto\text{\rm Prox}_{\lambda g}(x-\lambda\nabla f(x))$ is single-valued around $\bar{x}$. Then we deduce from Remark 5.3(i), the $\lambda$-criticality of $\bar{x}$, and (5.6) that $\nabla\varphi_{\lambda}(\bar{x})=0$. It follows from Remark 5.3(ii) that $\varphi_{\lambda}$ is of class $\mathcal{C}^{1,1}$ around $\bar{x}$, and thus we get that $\varphi_{\lambda}(x^{k})\to\varphi_{\lambda}(\bar{x})$ and $\|\nabla\varphi_{\lambda}(x^{k})\|\to 0$ as $k\to\infty$. Moreover, Lemma 5.5 implies that $\varphi_{\lambda}$ is strongly convex around $\bar{x}$, and hence there exists $\kappa>0$ such that
$\varphi_{\lambda}(x)\geq\varphi_{\lambda}(\bar{x})+\langle\nabla\varphi_{\lambda}(\bar{x}),x-\bar{x}\rangle+\frac{\kappa}{2}\|x-\bar{x}\|^{2}=\varphi_{\lambda}(\bar{x})+\frac{\kappa}{2}\|x-\bar{x}\|^{2},$